OpenStudy (anonymous):

state convergent or divergent if convergent find the sum justify all answers mathematically (n+1)/(2n+5) n=1 to infinity

OpenStudy (sirm3d):

use limit comparison test, with the DIVERGENT test series \[\frac{1}{n}\] or \[\frac{1}{2n}\]

OpenStudy (anonymous):

Would you be able to explain in more detail please?

OpenStudy (sirm3d):

sorry, i misread the question. you should use the divergence test, which states in \[\sum a_n\] if \[\lim_{n\rightarrow \infty} a_n \neq 0\] then the series is divergent.

OpenStudy (anonymous):

no worries, perfect then lol Thank you very much

OpenStudy (sirm3d):

yw